# Overcoming Reward Overoptimization via Adversarial Policy Optimization with Lightweight Uncertainty Estimation

### v1

@@ 5.1 AdvPO mitigates the over-optimization
Results. Figure 2a and Figure 2c illustrate how the golden reward (solid lines) and proxy reward (dashed lines) progress concerning policy optimization steps on Anthropic HH and TL;DR summarization datasets, respectively, while Figure 2b and Figure 2d capture the dynamics with respect to the square root KL divergence, i.e., $\sqrt{D_{\rm KL}(\pi_{\theta}||\pi_{\rm SFT})}$.
@@
Firstly, we note that PPO exhibits the issue of over-optimization across both datasets, characterized by a significant increase in proxy reward (blue dashed line), while the golden reward (blue solid line) begins to decline after reaching certain steps for both TL;DR and Anthropic datasets. However, by taking the estimated proxy reward conservatively, i.e. using AdvPO , we can mitigate over-optimization towards high but unreliable rewards, ensuring it stays within a reliable region (small KL divergence) with high golden rewards (red lines).
@@
Winnormal-$\uparrow$ Tie Losenormal-$\downarrow$ $\Delta$ Winnormal-$\uparrow$ Tie Losenormal-$\downarrow$ $\Delta$
LWUN-s PPO $33.5$ $39.5$ $27.0$uparrow 6.5 $70.3$ $3.00$ $26.7$uparrow 43.6
ENS-s $39.5$ $29.5$ $31.0$ normal-$\uparrow$ 8.5 $65.3$ $8.20$ $26.5$uparrow 38.8
AdvPO-noRef PPO $36.0$ $35.0$ $29.0$uparrow 7.0 $45.0$ $11.0$ $39.0$uparrow 6.0
LWUN-s $28.5$ $49.0$ $22.5$ normal-$\uparrow$ 6.0 $52.0$ $3.00$ $45.0$uparrow 7.0
AdvPO PPO $31.0$ $49.0$ $20.0$ normal-$\uparrow$ 11.0 $75.0$ $7.00$ $18.0$ normal-$\uparrow$ 57.0
PPO-ref $35.5$ $39.5$ $25.0$ normal-$\uparrow$ 10.0 $55.0$ $6.00$ $39.0$ normal-$\uparrow$ 16.0
@@
5.2 AdvPO results an improved policy
@@
Next, we tackle the question of whether AdvPO can effectively learn an improved policy in practical scenarios. We implement the standard RLHF pipeline on both Anthropic HH and TL;DR summarization datasets, where the reward model is trained based on human preferences. The algorithm’s performance is subsequently assessed by evaluating the quality of responses generated from the resulting policy.
Baselines. We compare AdvPO against the following:
  • $\bullet$
    PPO: The token-wise implementation of Proximal Policy Optimization (PPO) [32], treating each token as an action, that faces the over-optimization issue as it fully optimizes the policy based on the estimated reward, which may be inaccurate.
  • $\bullet$
    PPO-ref: A modified version of PPO which incorporates reference responses similar to that in Eq.(7).
  • $\bullet$
    ENS-s: An existing approach by [9, 11], which leverages an ensemble of reward models to estimate uncertainty and applies sample-wise uncertainty penalty to mitigate overoptimization. While our reward model is a LLama-7B, to roughly match the parameter size, we train three reward ensembles based on OpenLLaMA3B_v2 [13], as Meta AI has not officially released 3B models for LLaMA. OpenLLaMA [13] is a permissively licensed open-source reproduction of Meta AI’s LLaMA with comparable performance. The reward ensembles undergo fine-tuning on the same SFT data.
  • $\bullet$
    LWUN-s: This Light Weight UNcertainty approach utilizes reward uncertainty calculated through CI in Eq.(3). But it leverages sample-wise uncertainty penalization during policy optimization, akin to the methodology described in [9, 11].
  • $\bullet$
    AdvPO-noRef: To ablate our proposed AdvPO, we also add a variant of AdvPO which also conducts distributionally robust optimization centred around the reward model’s prediction confidence interval, however without incorporating reference responses.
Implementation Details. For both datasets, the reward model and policy model are initialized from LLama7B, fine-tuned using corresponding SFT data. Due to limited space, additional implementation details and hyperparameters, please refer to Appendix A.2.
Evaluation. While GPT-4 is commonly used as a proxy for human evaluation of generation quality [48, 40], we observed significant position bias issues in its output. In some cases, when flipping the position of two responses for the same pair of generations, GPT-4 yielded contradictory evaluation results.
Therefore, to get a fair assessment of the responses we use a combination of GPT-4 evaluation and human labelling as follows: For each paired response comparison, we query GPT-4 twice by swapping their positions. If GPT-4 consistently judges one answer as preferred to the other, we adopt GPT-4’s judgment. In cases where GPT-4 provides inconsistent judgments or declares a tie, we engage three individuals for annotations, and the majority vote among the manual annotations is considered the final evaluation. Given the expense and time-consuming nature of obtaining GPT4 and human annotations, we randomly select 200 prompts from the validation data of the Anthropic HH dataset and 100 prompts from the TL;DR dataset for evaluation. The GPT-4 evaluation prompt and human annotation instructions are detailed in App. B.
@@
Experiment Results. The results on Anthropic HH and TL;DR datasets are summarized in Table 1. We compare the models in pairs and report their win, lose, and tie ratios.
@@
We initially observe that, despite implementing sample-wise uncertainty penalization as described in [9, 11], leveraging lightweight-calculated uncertainty, as demonstrated by LWUN-s, aids in mitigating over-optimization during policy optimization. This results in an improved policy compared to PPO. Furthermore, LWUN-s outperforms ENS-s, highlighting the effectiveness of lightweight-calculated uncertainty compared to ensembles with similar parameter sizes.
@@
In Table 1, we see how our proposed AdvPO consistently outperforms all baselines, demonstrating particularly significant performance enhancements when the quality of reference responses is high. Specifically, on the TL;DR summarization dataset, where the reference responses exhibit considerable quality, AdvPO achieves substantial improvements. In contrast, the Anthropic HH dataset contains noise, with reference responses varying in quality, resulting in relatively smaller improvements.
Nevertheless, its advantage over PPO-ref still underscores the advantages of leveraging uncertainty conservatively to address overoptimization. Furthermore, in comparison to AdvPO-noRef, we observe that incorporating a reference leads to superior performance, preventing AdvPO from being overly conservative.
@@ 6 Related Work
Several recent works aim to directly learn the policy model without RL optimization, either through directly learning the policy model from preference data without involving reward modelling [21, 28, 1, 10, 15], or continuously mimicking responses ranked high by the reward model [10, 15, 45]. However, due to the inherent limitations of supervised learning, these approaches face challenges in generalization and are particularly vulnerable to out-of-preference data [20, 43].
There is also a line of work [9, 11, 46] that aims to directly address the overoptimization issue during policy optimization by penalizing samples with high uncertainty in the estimated reward. However, all of these approaches utilize the variance of reward ensembles with different seeds, either in pretraining or fine-tuning, as the measure of uncertainty. This not only incurs a high computational cost but also hinders achieving the maximum performance potential, given that the “scaling laws” generally advocate for larger reward models. Moreover, all of these methods adopt per-sample uncertainty penalization, which may be too conservative. Additionally, several theoretical works consider the accuracy of reward models in RLHF, primarily from an offline RL perspective [47, 49]. However, these works mainly contribute to the theoretical understanding of RLHF without any empirical experiments.
Adversarial Learning in RLHF. In addition to approaches countering over-optimization [9, 11, 46], recent work [7] proposes an adversarial optimization framework to update the reward and policy models iteratively. However, they utilize a min-max objective, where the inner optimization learns a policy to maximize rewards, while the outer minimization refines reward models based on provided gold preference data. Their inner optimization still directly relies on estimated rewards, thus suffering from the overoptimization problem. In contrast, our framework employs a max-min objective, where the inner minimization with a confidence region searches for rewards conservatively, based on which the policy is then maximized. Furthermore, their work is currently implemented only with rejection sampling as the LLM updating algorithm, unlike the RL optimization stage in our approach.
@@ 7 Conclusion and Future work
In this paper, we propose Adversarial Policy Optimization (AdvPO), a novel approach designed to tackle reward overoptimization in RLHF. First, we showcase the effectiveness of lightweight uncertainty quantification that relies solely on the last layer embeddings. Subsequently, we propose AdvPO  to tackle a distributionally robust optimization problem centred around the reward model’s prediction confidence interval for policy improvement. Empirical experiments conducted on the Anthropic HH and TL;DR summarization datasets demonstrate that AdvPO effectively mitigates the overoptimization issue without incurring the computational burden associated with ensembles. This results in improved policy in practical scenarios.
As our future work, we aim to explore constructing uncertainty estimation by considering layers beyond the last one. In our current experiments, we primarily focused on the last layer due to the abundance of literature [17, 31, 44, 5], and the absence of a principled way to select a middle layer without exhaustive search. Hence, future work would involve exploring whether considering different layers as well as combinations of them, can lead to more accurate uncertainty estimation and, consequently, the development of more advanced algorithms to address overoptimization. Besides, exploring different uncertainty estimations to actively select data for RLHF training could be a promising future direction for iterative improvement.
8 Impact Statement
This paper contributes to advancing the field of Machine Learning with a focus on enhancing the safety of Language Models (LMs). Our primary goal is to make LMs safer by actively downweighting unreliable prompt-response pairs in the policy optimization stage, ensuring that the LM aligns correctly with user preferences. In contrast to existing methods that lack this distinction, our approach aims to steer LMs towards greater safety for our society.
Appendix A Implementation Details
@@
A.1 Experimental details for synthetic setup in Section 5.1
@@
  • Supervised Fine-tuning. All reward models and policy models undergo fine-tuning from LLama7B [37] based on the Supervised Fine-tuning (SFT) data for each dataset. This aims to enhance instruction-following capabilities for the task. We set an initial learning rate of $5\times e^{-6}$, a batch size of $64$, and a context window length of $2048$ tokens.
@@
    Similar to the previous step, we train the reward models for up to three epochs and select the model that achieves the minimum loss on the validation dataset. The accuracy of the proxy reward models for Anthropic HH and TL;DR summarization datasets is 0.69 and 0.76, respectively.
    For both gold and proxy reward model training, we set the initial learning rate to $5e^{-6}$, a batch size of 32, and a context window length of 2048 tokens.
  • RL optimization: We apply both the standard PPO and the proposed AdvPO on the second half of the dataset for policy optimization. In both datasets, we split 90% for training and 10% for validation. For both algorithms, we train the model for 1500 steps, with an initial learning rate of $1e^{-6}$, a batch size of 64, and a context window length of 2048, consistent with previous procedures. For efficient online sampling, we set the maximum number of generated tokens to 512 and the KL coefficient $\beta$ to 0, following previous work [9]. For AdvPO, we use the chosen response for each prompt in the dataset as the reference response.
A.2 Experimental Details for Section 5.2
@@
In this section, we provide the GPT-4 prompts and human annotation instructions utilized for evaluating the TL;DR summarization and Anthropic HH datasets.
@@ Appendix C Theoretical Results
Theorem C.1.
The optimization problem in Eq.(7) is is equivalent to the following objective:
$\max_{\pi_{\theta}}\quad$ $\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}[r_{\phi}(x,y)-\frac{1}{\lambda^{*}}e(x,y)^{T}M^{-1}g]-\mathbb{E}_{x,y_{\rm ref}}[r_{\phi}(x,y_{\rm ref})-\frac{1}{\lambda^{*}}e(x,y_{\rm ref})^{T}M^{-1}g]$
[…]
where $e(x,y)$ denotes the last layer embedding of the prompt-response pair $x$ and $y$, and $g=\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[e(x,y)\right]-\mathbb{E}_{x,y_{\rm ref}}\left[e(x,y_{\rm ref})\right]$, and $\lambda^{*}=\sqrt{\frac{g^{T}M^{-1}g}{B}}$.
@@
With the definition of $e(x,y)$, the reward obtained under the projection weight $\phi$ is denoted as $r_{\phi}(x,y)=e(x,y)^{T}\phi$. The optimization problem in Eq.(7) can be rewritten as follows:

### v2

@@ 5.2 AdvPO mitigates reward overoptimization
Results. Figure 2a and Figure 2c illustrate how the golden reward (solid lines) and proxy reward (dashed lines) progress concerning policy optimization steps on both datasets, while Figure 2b and Figure 2d capture the dynamics with respect to the square root KL divergence, i.e., $\sqrt{D_{\rm KL}(\pi_{\theta}||\pi_{\rm SFT})}$.
@@
$\bullet$ PPO suffers from overoptimization, whereas AdvPO mitigates the issue. We can observe that PPO suffers from overoptimization across both datasets, characterized by a significant increase in proxy reward (blue dashed line), while the golden reward (blue solid line) begins to decline after reaching certain steps for both datasets. However, AdvPO mitigates over-optimization towards high but unreliable rewards, ensuring it stays within a reliable region (small KL divergence) with high golden rewards (red lines). Moreover, as shown in Figure 2 in the Appendix, the uncertainties of generated responses remain stable under AdvPO, unlike the significant increase observed with PPO. This again highlights the effectiveness of AdvPO in addressing over-optimization.
@@
[…]
LWUN-s PPO $33.5$ $39.5$ $27.0$uparrow 6.5 $50.0$ $20.0$ $30.0$uparrow 20.0
ENS-s $39.5$ $29.5$ $31.0$uparrow 8.5 $64.0$ $8.00$ $26.0$uparrow 38.0
AdvPO PPO $31.0$ $49.0$ $20.0$uparrow 11.0 $75.0$ $7.00$ $18.0$uparrow 57.0
PPO-ref $35.5$ $39.5$ $25.0$uparrow 10.0 $55.0$ $6.00$ $39.0$uparrow 16.0
@@
5.3 AdvPO results in an improved policy
@@
Next, we investigate whether AdvPO can effectively learn an improved policy in practical scenarios. Unlike the experimental setup described above, in this section, the RLHF pipeline is conducted by training the reward model based on human preferences using two datasets. The algorithm’s performance is then evaluated by assessing the quality of responses generated by the resulting policy.
Baselines. We compare AdvPO against the following: (1) PPO: the token-wise PPO algorithm [32]; (2)PPO-ref: a modified version of PPO which incorporates reference responses as in Eq.(6); (3) ENS-s: the ensemble-based approach to address over-optimization [9, 11] which quantifies uncertainty via ENS-3B as described in Section 5.1, utilizing three 3B reward ensembles. It then applies a sample-wise uncertainty penalty during the RL stage to counter overoptimization; (4) LoraEns [^1] We implement two versions: one with separate LoRA ensembles using different seeds, and another following the original paper. For each dataset, we report the better result of the two.: a recent work [45] that trains LoRA-based reward ensembles to save memory costs while using sample-wise uncertainty penalties during the RL stage. Five LoRA ensembles are trained, with LoRA dimensions set at 32 and LoRA-alpha at 64; (5) LWUN-s: the approach that utilizes reward uncertainty calculated through CI, but through sample-wise uncertainty penalization during the PPO stage; (6) AdvPO-noRef: a variant of AdvPO without incorporate reference responses in Eq. (6).
Implementation & Evaluation Details. While GPT-4 is often employed to gauge generation quality, we noted significant position bias issues in its output. Thus, for a fair assessment of responses, we combine GPT-4 evaluation with human labelling. For additional implementation and evaluation details, please refer to Appendix A.4 and B.
@@
Results. We compare the models in pairs and report their win/lose/tie ratios in Table 1.
@@
$\bullet$ Lightweight uncertainty works even with sample-wise penalization. Despite implementing sample-wise uncertainty penalization [9, 11], leveraging lightweight-calculated uncertainty, as demonstrated by LWUN-s, aids in mitigating overoptimization during policy optimization. This results in an improved policy compared to PPO. Furthermore, LWUN-s outperform ENS-s, highlighting the effectiveness of lightweight uncertainty compared to ensembles with similar parameter sizes.
@@
$\bullet$ AdvPO outperforms all baselines, with high-quality reference responses further enhancing its performance. From Table 1, it’s evident that AdvPO consistently outperforms all baselines, showing significant performance improvements, especially when the quality of reference responses is high. Specifically, on the TL;DR dataset, where the reference responses exhibit considerable quality, AdvPO achieves substantial improvements. In contrast, the Anthropic HH dataset contains noise, with reference responses varying in quality, resulting in relatively smaller improvements. Still, its advantage over PPO-ref highlights the benefits of conservatively leveraging uncertainty to address overoptimization. Additionally, compared to AdvPO-noRef, incorporating a reference improves performance, ensuring AdvPO isn’t overly conservative.
@@ 6 Related Work
Several recent works [22, 28, 1, 10, 15] aim to directly learn the policy model without RL optimization. However, due to supervised learning’s inherent limitations, these approaches encounter challenges in generalization and are especially susceptible to out-of-preference data [21, 42]. Another line of work [9, 11, 45] aims to directly address the overoptimization issue during policy optimization by penalizing samples with high reward uncertainty, measured as variance among reward ensembles. However, fully-finetuned ensembles [9, 11] not only incur high computational costs but also hinder achieving maximum performance, as the ”scaling laws” generally advocate for larger reward models. On the other hand, while LoRA-based ensembles [45] reduce memory requirements, they still incur additional training costs and computational overhead due to querying each ensemble for each sample to calculate reward and uncertainty. Additionally, several theoretical works consider the accuracy of reward models in RLHF, primarily from an offline RL perspective [46, 48]. However, these works mainly contribute to the theoretical understanding of RLHF without any empirical experiments.
Adversarial Learning in RLHF. In addition to approaches countering over-optimization [9, 11, 45], recent work [7] proposes an adversarial optimization framework for iteratively updating reward and policy models. However, they utilize a min-max objective, where the inner optimization learns a policy to maximize rewards, while the outer minimization refines reward models based on provided gold preference data. Their inner optimization still directly relies on estimated rewards, thus suffering from the overoptimization problem. In contrast, our framework employs a max-min objective, where the inner minimization with a confidence region searches for rewards pessimistically, based on which the policy is then maximized. Furthermore, their work is currently implemented only with rejection sampling as the LLM updating algorithm, unlike the RL optimization stage in our approach.
@@ 7 Conclusion and Future work
In this paper, we introduce AdvPO, a novel approach designed to address reward overoptimization in RLHF, motivated by the effectiveness of our proposed lightweight uncertainty quantification method. Empirical experiments on the Anthropic HH and TL;DR datasets show that AdvPO effectively mitigates overoptimization without the computational burden of ensembles, leading to improved policy in practical scenarios.
Limitations: In this work, we only considered constructing uncertainty from the last layer of the reward model. Future work could consider constructing possibly more accurate estimates with intermediate layers as well. In addition, we only explored the use of uncertainty for model training. Exploring uncertainty estimations to actively select data for RLHF training could be a promising future direction for iterative improvement. Lastly, additional experiments on larger scale model i.e in the order of 70B, would be interesting, however this is outside the scope of this paper.
Appendix A Experimental details
A.1 Datasets.
We utilized the following two widely adopted datasets for RLHF to carry out our empirical investigation.
  • $\bullet$
  • $\bullet$
A.2 Implementation Details
@@
Choice of $b$ in AdvPO. As shown in the proof of Theorem 3.1, we have $||\phi^{*}-\hat{\phi}||_{M_{D}}\leq b$ holds with probability $1-\delta$. The term $b$ is closely related to $\delta$, where the smaller $\delta$ is (i.e., the larger probability the confidence region holds), the larger $b$ is.
The choice of $\delta$ significantly impacts algorithm performance. If $\delta$ is too small, even though the resulting larger confidence ball from the larger $b$ will likely cover the ground-truth projection weight $\phi^{*}$ and thus the ground-truth reward, taking pessimistic reward within such a larger confidence region can lead to overly pessimistic PPO learning, resulting in worse performance. Conversely, if $\delta$ is too large, the ground-truth reward may not be covered by the resulting confidence region, leading to incorrect optimization direction for PPO.
It is challenging to pre-determine the optimal $\delta$ for a dataset. Therefore, following previous work [43], in our experiments, we take $||\phi^{*}-\hat{\phi}||^{2}_{M_{D}}\leq b^{2}=B$ and treat $B$ as a hyperparameter.
Hyperparameter Tuning. A good $B$ should: (1) prevent over-optimization by searching for reward predictions within the resulting confidence region that are most pessimistic about the current policy, guiding policy optimization (PPO); (2) avoid excessive pessimism to ensure a well-performing policy.
In our experiment, we monitor the average uncertainty in each batch during policy optimization and observe how it changes over time to inspect the first point. For the second point, every 100 optimization steps in each run, we utilize the current policy to generate responses for prompts in the validation dataset and record the average reward on the validation dataset. The checkpoint that achieves the highest reward in the validation dataset during training is selected as the resulting model for this run. We then select the best $B$ from the list $[1,5,10,15]$, which stabilizes uncertainty in the later stages of optimization while the resulting model achieves the highest reward on the validation dataset.
A.3 Experimental details for synthetic setup in Section 5.2
All our experiments were run on a cluster of 8xA100 GPUs with 100 CPUs and 100 GB RAM. Reward modelling took roughly on average 1 day whereas PPO took roughly 2 days.
@@
  • Supervised Fine-tuning. All reward models and policy models undergo fine-tuning from LLama7B [36] based on the Supervised Fine-tuning (SFT) data for each dataset. This aims to enhance instruction-following capabilities for the task. We set the learning rate to $5e^{-6}$ for the Anthropic HH dataset and $3e^{-5}$ for the TL;DR dataset. In both cases, the batch size is 64, and the models are trained for 5 epochs. The checkpoint with the lowest loss on the validation dataset is selected.
@@
    Similar to the previous step, we train the reward models for up to three epochs and select the model that achieves the minimum loss on the validation dataset. The accuracy of the proxy reward models for the Anthropic HH and TL;DR summarization datasets on the validation datasets is 0.69 and 0.76, respectively. For both gold and proxy reward model training, we set the initial learning rate to $5e^{-6}$, a batch size of 64, and a context window length of 2048 tokens.
  • RL optimization: We apply both the standard PPO and the proposed AdvPO on the second half of the dataset for policy optimization. In both datasets, we split 90% for training and 10% for validation. For both algorithms, we train the model for 1500 steps, with an initial learning rate of $1e^{-6}$, a batch size of 64, and a context window length of 2048, a PPO value clip threshold of 0.2, consistent with previous procedures. For efficient online sampling, we set the maximum number of generated tokens to 512 and the KL coefficient $\beta$ to 0 to encourage the most severe over-optimization scenario, following previous work [9]. For AdvPO, we use the chosen response for each prompt in the dataset as the reference response.
    In each single run, every 100 optimization steps, we use the current policy to generate responses for prompts in the validation dataset and record the average reward on the validation dataset. We then select the checkpoint that achieves the highest reward in the validation dataset during training as the resulting model for this run.
A.4 Experimental Details for Section 5.3
@@
While GPT-4 is commonly used as a proxy for human evaluation of generation quality [47, 39], we observed significant position bias issues in its output. In some cases, when flipping the position of two responses for the same pair of generations, GPT-4 yielded contradictory evaluation results. Therefore, to get a fair assessment of the responses we use a combination of GPT-4 evaluation and human labelling as follows: For each paired response comparison, we query GPT-4 twice by swapping their positions. If GPT-4 consistently judges one answer as preferred to the other, we adopt GPT-4’s judgment. In cases where GPT-4 provides inconsistent judgments or declares a tie, we engage three individuals for annotations, and the majority vote among the manual annotations is considered the final evaluation. Given the expense and time-consuming nature of obtaining GPT4 and human annotations, we randomly select 200 prompts from the validation data of the Anthropic HH dataset and 100 prompts from the TL;DR dataset for evaluation.
The GPT-4 prompts and human annotation instructions used for evaluating the TL;DR summarization and Anthropic HH datasets are as follows. There was no risk to annotators.
@@ Appendix C Theoretical Results
Proof of Theorem 3.1:
@@
For ease of illustration, we denote the parameters in $e(x,y)$ as $\hat{w}$, thus $\hat{\varphi}=(\hat{\phi},\hat{w})$. Additionally, We also use $\phi^{*}$ and $w^{*}$ to denote the unknown ground-truth of $\hat{\phi}$ and $\hat{w}$ respectively. At a high level, the derivation of the reward uncertainty $|r^{*}(x,y)-r_{\hat{\varphi}}(x,y)|$ consists the following steps:
  • Step 1: Obtain the confidence region of the learned projection weight $\hat{\phi}$ with preference dataset $D$, such that with probability $1-\delta$, $\|\phi^{*}-\hat{\phi}\|_{M_{D}}\leq b$ holds, where $b$ is a term related to $D$ and $\delta$. Typically, the smaller $\delta$ is, the larger $b$ is.
  • Step 2: Derive the reward uncertainty $|r^{*}(x,y)-r_{\hat{\varphi}}(x,y)|$ based on the confidence region of $\hat{\phi}$.
Proof of Step 2. We first elaborate how to derive the reward uncertainty $|r^{*}(x,y)-r_{\hat{\varphi}}(x,y)|$ given $\|\phi^{*}-\hat{\phi}\|_{M_{D}}\leq b$.
Under the assumption of infinitely wide networks and a positive definite neural tangent kernel matrix, we first demonstrate $r^{*}(x,y)$ can be approximated by a linear function of $e(x,y)$ and $\phi^{*}$ along with terms related to init parameters points $(\phi_{0},w_{0})$, i.e.:
$r^{*}(x,y)=e(x,y)^{T}\phi^{*}+\phi_{0}^{T}\cdot\nabla_{w}e(x,y;w_{0})\cdot(w^{*}-w_{0}),$
where $e(x,y;w_{0})$ is the last layer embedding generated based on the initial parameters $(\phi_{0},w_{0})$. The proof closely resembles the proof of Lemma A.1 in [43], with the distinction that the convergence of the neural tangent kernel is not exclusive to fully-connected architectures; it can be extended to transformers, as demonstrated in recent work [44, 24].
While $\|w^{*}-w_{0}\|$ is bounded as in Lemma A.1 in [43], and $\nabla_{w}e(x,y;w_{0})$ is the derivative with respect initial parameter $w_{0}$. If $w_{0}$ are drawn from standard Gaussians (i.e. in the NTK parametrization), as widths tend to infinity, $\nabla_{w}e(x,y;w_{0})$ is bounded as shown by Lemma A.2 in [43]. Consequently, with probability $1-\delta$, we have
$|r^{*}(x,y)-r_{\hat{\varphi}}(x,y)|$ $=|e(x,y)^{T}\phi^{*}-e(x,y)^{T}\hat{\phi}+\phi_{0}^{T}\cdot\nabla_{w}e(x,y;w_{0})\cdot(w^{*}-w_{0})|$
$\leq|e(x,y)^{T}\phi^{*}-e(x,y)^{T}\hat{\phi}|+\text{const}$ (8)
Utilizing the inequality $|u^{T}v|\leq\|u\|_{A}\|v\|_{A^{-1}}$ for any postive definite matrix $A$, derived from Cauchy-Schwarz inequality, we further obtain:
$|e(x,y)^{T}\phi^{*}-e(x,y)^{T}\hat{\phi}|\leq\|e(x,y)\|_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{D}^{-1}}\|\phi^{*}-\hat{\phi}\|_{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}M_{D}}$ (9)
Proof of Step 1. Next, we present a proof for step 1, deriving the confidence region of $\hat{\phi}$. From the reward modeling, we have:
$\hat{\phi}=\operatorname*{arg\,min}_{\phi}L_{D}=\sum_{(x,y_{c},y_{r})\in D}\sum_{k\in\{c,r\}}i_{x,k}L_{x,k}=\sum_{(x,y_{c},y_{r})\in D}\sum_{k\in\{c,r\}}i_{x,k}\log\left(\frac{\exp(e(x,y_{k})^{T}\phi)}{\exp(e(x,y_{c})^{T}\phi)+\exp(e(x,y_{r})^{T}\phi)}\right)$ (10)
where $i_{x,c}=1$ and $i_{x,r}=0$. Let $p_{x}^{k}:=\frac{\exp(e(x,y)^{T}\hat{\phi})}{\exp(e(x,y_{c})^{T}\hat{\phi})+\exp(e(x,y_{r})^{T}\hat{\phi})},k\in\{c,r\}$. Since $\hat{\phi}$ is the minimizer regarding $L_{D}$, setting the derivative of $L_{D}$ with respect to $\hat{\phi}$ as zero, we have :
$\sum_{(x,y_{c},y_{r})\in D}\sum_{k\in\{c,r\}}(p_{x,k}-i_{x,k})e(x,y_{k})=0$
Denote
$\mu(\phi,e(x,y_{k}))=\frac{\exp(e(x,y_{k})^{T}\phi)}{\sum_{k\in\{c,r\}}\exp(e(x,y_{k})^{T}\phi)}$
and
$G(\phi)=\sum_{D}\sum_{k\in\{c,r\}}[\mu(\phi,e(x,y_{k}))-\mu(\phi^{*},e(x,y_{k}))]e(x,y_{k}).$
Following the Step 1 of Theorem 1 in [20], we can derive :
$\|\hat{\phi}-\phi^{*}\|^{2}_{M_{D}}\leq\frac{1}{\kappa^{2}}\|G(\hat{\phi})\|_{M_{D}^{-1}}^{2},$
where $\kappa:=\inf_{\|\phi^{*}-\phi\|\leq 1}\dot{\mu}(\phi,e(x,y))\geq 0,\forall e(x,y)$ (Assumption 1 in [20]). Then Lemma 3 in [20] further bounds $\|G(\hat{\phi})\|_{M_{D}^{-1}}^{2}$ by a term related to D and $\delta$, resulting $\|\phi^{*}-\phi\|_{M_{D}}\leq b$ holding with probability $1-\delta$.
Thus we conclude the proof.
Lemma C.1.
The inclusion of reference responses prevents AdvPO from being overly or wrongly pessimistic by enforcing policy optimization towards the direction of the reference responses while optimizing against pessimistic rewards.
Proof.
Let $\hat{\phi}_{\rm ref}^{*}$ and $\hat{\phi}_{\rm noref}^{*}$ represent the derived projection weights of the inner optimization of the max-min objective in Eq.(7) with or without reference responses, respectively. Denote $g_{\rm\pi_{\theta}}=\mathbb{E}_{x,y\sim\pi_{\theta}(\cdot|x)}\left[e(x,y)\right]$ and $z_{\rm ref}=\mathbb{E}_{x,y_{\rm ref}}\left[e(x,y_{\rm ref})\right]$ Thus the policy optimization objective for max-min objective ¡em¿with¡/em¿ reference responses (i.e., Eq.(7)) is
$J_{\rm ref}=\max_{\pi_{\theta}}\quad g_{\rm\pi_{\theta}}^{T}\hat{\phi}_{\rm ref}^{*}-z_{\rm ref}^{T}\hat{\phi}_{\rm ref}^{*}=\max_{\pi_{\theta}}\quad g_{\rm\pi_{\theta}}^{T}\hat{\phi}_{\rm ref}^{*}.$
The last equality holds because the second term is a constant given $\hat{\phi}_{\rm ref}^{*}$, thus subtracting it will not affect the resulted optimal policy. Similarly, we can derive the policy optimization objective for max-min objective withoutreference responses:
$J_{\rm noref}=\max_{\pi_{\theta}}\quad g_{\rm\pi_{\theta}}^{T}\hat{\phi}_{\rm noref}^{*}.$
Following a similar procedure as proof of Theorem4.1 and replacing $g$ in Eq.(14) by $g_{\rm\pi_{\theta}}-z_{\rm ref}$ and $g_{\rm\pi_{\theta}}$, we can derive the closed-form solution of $\hat{\phi}_{\rm ref}^{*}$ and $\hat{\phi}_{\rm noref}^{*}$. By plugging them into $J_{\rm ref}$ and $J_{\rm noref}$, we can get:
$J_{\rm ref}=\max_{\pi_{\theta}}\quad g_{\rm\pi_{\theta}}^{T}\hat{\phi}_{\rm ref}^{*}=g_{\rm\pi_{\theta}}^{T}\hat{\phi}-\frac{1}{\lambda^{*}_{\rm ref}}g_{\rm\pi_{\theta}}^{T}M_{D}^{-1}g_{\rm\pi_{\theta}}+\underbrace{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{1}{\lambda^{*}_{\rm ref}}g_{\rm\pi_{\theta}}^{T}M_{D}^{-1}z_{\rm ref}}_{(A_{\rm ref})},$
and
$J_{\rm noref}=\max_{\pi_{\theta}}\quad g_{\rm\pi_{\theta}}^{T}\hat{\phi}_{\rm noref}^{*}=g_{\rm\pi_{\theta}}^{T}\hat{\phi}-\frac{1}{\lambda^{*}_{\rm noref}}g_{\rm\pi_{\theta}}^{T}M_{D}^{-1}g_{\rm\pi_{\theta}}$
where $\lambda^{*}_{\rm ref}$ and $\lambda^{*}_{\rm noref}$ are Lagrangian multipliers derived from the optimization process.
We can observe that both $J_{\rm ref}$ and $J_{\rm noref}$ aim to prevent the policy from moving in the direction of high uncertainty by minimizing $g_{\rm\pi_{\theta}}^{T}M_{D}^{-1}g_{\rm\pi_{\theta}}$. However, $J_{\rm ref}$ includes an additional term $A_{\rm ref}$ compared to $J_{\rm noref}$. This term encourages the policy $\pi_{\theta}$ to move towards the reference responses $z_{\rm ref}=\mathbb{E}_{x,y_{\rm ref}}[e(x,y_{\rm ref})]$. With reasonably good reference responses, i.e., $z_{\rm ref}^{T}\phi^{*}>0$, the additional term $A_{\rm ref}$ guides the policy in a more accurate optimization direction, preventing AdvPO from being overly or wrongly pessimistic. ∎
Proof of Theorem 4.1:
Proof.
@@
With the definition of $e(x,y)$, the reward obtained under the projection weight $\phi$ is denoted as $r_{\phi}(x,y)=e(x,y)^{T}\phi$. With $B=b^{2}$, the optimization problem in Eq.(6) can be rewritten as follows: